At 230K temp,solid Ar Gas has a density of 1.65 gm/cc. Ar has a spherical geometric shape and it’s radius is 1.54 x 10^-8 cm. How much percentage of void is there in solid Ar?

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii Rashid
Volume of one molecule = 4/3 ( 3.14) (1.54*10-8)
= 1.53*10-23
Volume of molecules of 1.65g Ar = 1.65/40 * N0 *1.53* 10-23 = 0.380 cm3
Volume of solid content of 1.65g = 1cm3
empty space = 1-0.380 = 0.620
% of empty space = 62%